600 waves pass through a point on the surface of a pond in 1 minute. the distance between two consecutive crests is 35cm. find the velocity of waves
 

Grade:10

1 Answers

Mahima Kanawat
6 years ago
Dear student ...velocity of wave = nλ. Frequency(n) is the number of waves per second , Here n=10 and λ = 35×10-² m , so v=10×35×10-²=3.5 m/s .....With Regards ....Mahima...(AskIITians Forum Expert )